Evaluate the integral of (e^(√x)/√x) dx from 1 to 4.


1
Expert's answer
2021-03-01T06:36:42-0500

Evaluate "\int" (e^("\sqrt{x}" )/"\sqrt{x}" ) dx from 1 to 4

Solution

let v = "\sqrt{x}" ; hence dv = d/dx ("\sqrt{x}" ) = 1/(2"\sqrt{x}" ) dx

So "\int" (e^("\sqrt{x}" )/"\sqrt{x}" ) dx = "\int" 2ev dv

Applying constant multiple rule: "\int" c f(v) dv = c"\int" f(v) dv with c = 2 and f(v) = ev

"\int" 2ev dv = 2"\int" ev dv

The integral of the exponential function is: "\int" ev dv = ev

Therefore: 2"\int" ev dv = 2ev

Recall that v = "\sqrt{x}"

Therefore "\int" (e^("\sqrt{x}" )/"\sqrt{x}" ) dx = 2e^("\sqrt{x}" )

Applying the limits from 1 to 4 to the integral result we get:

2e^("\sqrt{4}" ) - 2e^"(\sqrt{1})" = 9.341548541

Answer: 9.341548541